The legs of a right triangle are 15 meters and 9 meters. Find the length of the hypotenuse?
Answer by nerdybill(7384)

The legs of a right triangle are 15 meters and 9 meters. Find the length of the hypotenuse?
applying the Pythagorean theorem:
h^2 = 15^2 + 9^2
h^2 = 225 + 81
h^2 = 306
h = 17.49 meters